OpenStudy (anonymous):

What is the derivative of: x-1/x^(1/2) Their answer is 1/2x^(1/2)+1/2x^(3/2)

zepdrix (zepdrix):

Simplify before taking a derivative, so as to avoid the quotient rule :) Split it into 2 fractions.\[\huge \frac{x}{x^{1/2}}-\frac{1}{x^{1/2}}\]

zepdrix (zepdrix):

Simplifying further gives us... \[\huge x^{1/2}-x^{-1/2}\]If this step confuses you, let me know :D

zepdrix (zepdrix):

Taking the derivative, applying the power rule to each term gives us... \[\huge \frac{1}{2}x^{-1/2}-\left(-\frac{1}{2}x^{-3/2}\right)\]

OpenStudy (anonymous):

Yes Thats it! Thanks splitting into 2 fractions :/ forgot about that step